In the first-order reaction A→products, it is found that 99 % of the original amount of reactant A decomposes in 144 min .
Answer #1

A° = 100

A(final) = 100-99 = 1

Time = 144 min

Ln( A°/A ) = rate constant * time

Ln(100/1) = rate constant * 144

Rate constant = 0.032

Half Life = 0.693 / rate constant

= 0.693 / 0.032

= 21.67 min

If you have any questions please comment

If you satisfied with the solution please rate it thanks